OpenStudy (anonymous):

Simplify x+1/2x^2+5x+3 divided by 2/2x+3 Need To Show All Work Will Give A medal To the Best Responce

OpenStudy (campbell_st):

well factor the denominator of the 1st fraction... then remember than when you divide by a fraction you need to flip it and multiply so its \[\frac{x +1}{2x^2 + 5x + 3} \times \frac{2x +3}{2}\] after factoring the denominator of the 1st fraction you'll see common factors that cancel. good luck